# Attention-based Shape-Deformation Networks for Artifact-Free Geometry Reconstruction of Lumbar Spine from MR Images

### v2

@@ Abstract
Lumbar disc degeneration, a progressive structural wear and tear of lumbar intervertebral disc, is regarded as an essential role on low back pain, a significant global health concern. Automated lumbar spine geometry reconstruction from MR images will enable fast measurement of medical parameters to evaluate the lumbar status, in order to determine a suitable treatment. Existing image segmentation-based techniques often generate erroneous segments or unstructured point clouds, unsuitable for medical parameter measurement. In this work, we present TransDeformer: a novel attention-based deep learning approach that reconstructs the geometry of the lumbar spine with high spatial accuracy and mesh correspondence across patients, and we also present a variant of TransDeformer for error estimation. Specially, we devise new attention modules with a new attention formula, which integrate image features and tokenized contour features to predict the displacements of the points on a shape template without the need for image segmentation. The deformed template reveals the lumbar spine geometry in an image. Experiment results show that our TransDeformer generates artifact-free geometry outputs, and its variant predicts the error of a reconstructed geometry. Our code is available at https://github.com/linchenq/TransDeformer-Mesh.
@@ 2 Related Works
Graph convolution network (GCN) has demonstrated ground-breaking performance in geometry modeling by leveraging the structural relationships among vertices and edges [20]. GCNs can predict displacements of points/vertices of a template using image features. Wickramasinghe et al. proposed Voxel2Mesh, a GCN-based model with adaptive sampling and pooling strategies for geometry reconstruction of liver, hippocampus and synaptic junction [21]. Although this Voxel2Mesh approach outperformed several segmentation-based methods, it generated an uncertain number of vertices, thus lacking mesh correspondence across patients. Kong et al. proposed a compound framework fusing a UNet with GCN to deform templates for whole-heart geometry reconstruction on cardiac datasets [22]. As will be shown in our experiments, such a UNet-GCN approach still generates geometry artifacts in our application.
Instead of treating a geometry as a graph, some studies regards geometry reconstruction as a template-image registration task, wherein a diffeomorphic displacement field [23] is generated by a network to transform an initial template mesh to the target. For example, Pak et al. proposed DeepCarve that predicted a regular-grid displacement field by a UNet and used it to deform a pre-defined template of left-ventricle and aorta complex on a dataset of 80 cardiac CT scans [24]. As will be shown in our experiments, such a UNet-Disp approach is very sensitive to template initialization, such that some perturbation on initial position of the template may aggravate the quality of the output geometry.
@@ 3 Methods
We designed two novel networks for lumbar spine geometry reconstruction, named UNet-DeformSA and TransDeformer. The UNet-DeformSA model (Fig. 2) has a UNet backbone and a shape self-attention (SSA) mechanism that sets it apart from the existing models. The TransDeformer model (Fig. 3) gets rid of the UNet backbone by using cross-attention between shape and image as well as shape self-attention and image self-attention, which further improves the performance. The attention mechanisms in ViT and its successors [26] assume that tokens are located on a regular grid and therefor could not be used in our networks. To enable the new attention modules in our networks, we developed new equations for attention with relative position embedding. Furthermore, we modified TransDeformer for error estimation, i.e., to estimate the error of a reconstructed geometry. In the following sections, we provide the details of our models. For information not revealed due to the 10-page limit, we refer the reader to the source code.
@@ 3.1 Attention with Relative Position Embedding
To better explain Eq.(2), we describe this attention from the perspective of individual tokens. We incorporate the sinusoidal functions to capture the relative position information, which is illustrated as:
@@
The final output of an attention layer is the addition of a term similar to Eq.(1) and a term related to the relative positions. The $i$-th row of the output is calculated by
@@ 3.3.1 Image Self-Attention (ISA) Module
Figure 4: The overview of image self-attention (ISA) module.
@@ 3.3.3 Shape Self-Attention (SSA) Module
The architecture is depicted in Fig. 5. The SSA module aims to learn contextual dependencies across the points of a shape, where individual points are associated with embedding vectors that collectively form the shape feature matrix ${X_{S}}\in{\mathbb{R}^{N_{p}\times E}}$. $N_{p}$ is the number of points/tokens of the shape and $E$ is the embedding/feature dimension. A row in the matrix ${X_{S}}$ is a token in the attention mechanism described in Section A. For self-attention calculation, the position of a token is defined on the undeformed template with normalized coordinates between -1 and 1. The SSA module has multiple self-attention layers to generate a sequence of feature matrices.
@@ 3.5 Error Estimation Using a Modified TransDeformer
Figure 7: The overview of the shape error estimation mode (unit: mm).
@@ 4.1 Dataset
Our dataset [11] consists of de-identified lumbar spine MR images of 100 patients. Three experts annotated and reviewed each mid-sagittal MR image to ensure accuracy and consistency. Each mid-sagittal MR image was manually marked with boundaries and landmarks of the lumbar vertebrae and discs, which is guided by the established protocol [7]. The images are resized to 512 $\times$ 512. Each image is accompanied with 12 segmentation masks, corresponding to the background, 6 vertebrae (named L1, L2, L3, L3, L4, L5, and S1), and 5 discs (named D1, D2, D3, D4, and D5). We split the entire dataset into 70 training samples, 10 validation samples, and 20 test samples. A special data augmentation method was applied to the original dataset, resulting in 7000 training samples, 250 validation samples, and 2500 test samples. The data augmentation method is described in details in our technical report [11]. Our experiments were conducted on the augmented datasets that are publicly available [11]. Data diversity is illustrated in Fig. 8 by comparing the template with the ground-truth meshes, for which the centroids are aligned.
Figure 8: Distribution of Averaged Point-to-Point Distance (APPD) between the template mesh and the ground truth meshes of the patients in the test set. APPD is defined in Section 4.4.
4.2 Comparison with Related Approaches
4.2.1 UNet-GCN Model
This model was originally proposed by Kong et al. [22] for reconstructing human heart geometries from 3D volumetric images. The model has a UNet backbone pretrained for binary segmentation, and the encoder of the UNet is used as the image feature extractor at different scales. Three GCN modules are used to deform a template sequentially by using image features sampled at the points of the template. For our application, we followed the original design as close as possible, and made necessary adjustment by changing the 3D UNet to a 2D UNet to handle 2D sagittal images of lumbar spline. The UNet-GCN model uses GCN layers to fuse information from neighbour points for displacement prediction. Our UNet-DeformSA model has shape self-attention modules to integrate information from different points that could be far away from each other (i.e., utilizing long-range dependency), which is the key difference from and advantage over the UNet-GCN model.
4.2.2 UNet-Disp Model
4.2.3 UNet-MLP Model
4.3 Implementation
We used Pytorch ver. 2.0.0 [28] to implement the overall frameworks, and used Pytorch Geometric ver. 2.3.0 [29] to implement GCNs. We used the Adam optimizer with a fixed learning rate of 1e-4 and a batch size of 20. The training of each model completed within 24 hours on a single NVIDIA RTX A6000 with 48GB VRAM.
4.4 Metrics and Results of Geometry Reconstruction
To measure the accuracy of a reconstructed lumbar object (a vertebra or a disc), we assessed two key metrics: (1) Averaged Point-to-Point Distance (APPD) between a predicted shape and the corresponding ground-truth shape and (2) Dice Similarity Coefficient (Dice), measuring the overlap between areas enclosed by predicted and ground-truth shapes. For each metric, mean, standard deviation (std), the worse case, the $95$th or $5$th percentile among the test samples are calculated.
@@ 4.8 Hyperparameter Tuning on Validation Set
For the UNet-DeformSA model, we investigated the effect of the number of attention layers in the SSA module, and the results on the validation set are reported in Table 8. Based on this results, 2 layers of attention are used in the main experiments. For the UNet-GCN model, we evaluated different types of GCN layers and chose ResGatedGraphConv [30] based on the results on the validation set. For the UNet-Disp model, we explored the hyperparameter $\sigma$ and set it to 1 based on the results on the validation set. The backbone UNet structure is the same for the three models.
@@
Node: TransDeformer results (lower is better)
@@
Node: UNet-DeformSA results (lower is better)
@@ 5 Discussion
All of the current ML models for geometry reconstruction, including our models, are data-drive and therefore have no guarantee in output accuracy for an input image. Thus, the geometry output from a ML model has to be checked (and modified if necessary) by a human operator to ensure high accuracy for clinical use. The shape error estimation model, which is a modified version of TransDeformer, facilitates this quality control process by ranking the model-reconstructed geometries by the estimated errors. To reduce human negligence, those with larger (estimated) errors will be checked first when the operator is vigilant; and those with smaller (estimated) errors will be checked later when the operator possibly becomes fatigue.

### v3

@@ Abstract
Lumbar disc degeneration, a progressive structural wear and tear of lumbar intervertebral disc, is regarded as an essential role on low back pain, a significant global health concern. Automated lumbar spine geometry reconstruction from MR images will enable fast measurement of medical parameters to evaluate the lumbar status, in order to determine a suitable treatment. Existing image segmentation-based techniques often generate erroneous segments or unstructured point clouds, unsuitable for medical parameter measurement. In this work, we present UNet-DeformSA and TransDeformer: novel attention-based deep neural networks that reconstruct the geometry of the lumbar spine with high spatial accuracy and mesh correspondence across patients, and we also present a variant of TransDeformer for error estimation. Specially, we devise new attention modules with a new attention formula, which integrate image features and tokenized contour features to predict the displacements of the points on a shape template without the need for image segmentation. The deformed template reveals the lumbar spine geometry in an image. Experiment results show that our networks generate artifact-free geometry outputs, and the variant of TransDeformer can predict the errors of a reconstructed geometry. Our code is available at https://github.com/linchenq/TransDeformer-Mesh.
@@ 2 Related Works
Graph convolution network (GCN) has demonstrated ground-breaking performance in geometry modeling by leveraging the structural relationships among vertices and edges [20]. GCNs can predict displacements of points/vertices of a template using image features. Wickramasinghe et al. proposed Voxel2Mesh, a GCN-based model with adaptive sampling and pooling strategies for geometry reconstruction of liver, hippocampus, and synaptic junction [21]. Although this Voxel2Mesh approach outperformed several segmentation-based methods, it generated an uncertain number of vertices, thus lacking mesh correspondence across patients. Kong et al. proposed a compound framework fusing a UNet with GCN to deform templates for whole-heart geometry reconstruction on cardiac datasets [22]. As will be shown in our experiments, such a UNet-GCN approach still generates geometry artifacts in our application.
Instead of treating a geometry as a graph, some studies regard geometry reconstruction as a template-image registration task, wherein a diffeomorphic displacement field [23] is generated by a network to transform an initial template mesh to the target. For example, Pak et al. proposed DeepCarve that predicted a regular-grid displacement field by a UNet and used it to deform a pre-defined template of left-ventricle and aorta complex on a dataset of 80 cardiac CT scans [24]. As will be shown in our experiments, such a UNet-Disp approach is very sensitive to template initialization, such that some perturbation on initial position of the template may aggravate the quality of the output geometry.
@@ 3 Methods
We designed two novel networks for lumbar spine geometry reconstruction, named UNet-DeformSA and TransDeformer. The UNet-DeformSA model (Fig. 2) has a UNet backbone and a shape self-attention (SSA) mechanism that sets it apart from the existing models. The TransDeformer model (Fig. 3) gets rid of the UNet backbone by using cross-attention between shape and image as well as shape self-attention and image self-attention, which further improves the performance. The attention mechanisms in ViT and its successors [26] assume that tokens are located on a regular grid and therefore could not be used in our networks. To enable the new attention modules in our networks, we developed new equations for attention with relative position embedding. Furthermore, we modified TransDeformer for error estimation, i.e., to estimate the error of a reconstructed geometry. In the following sections, we provide the details of our models. For information not revealed due to the 10-page limit, we refer the reader to the source code.
@@ 3.1 Attention with Relative Position Embedding
To better explain (2), we describe this attention from the perspective of individual tokens. We incorporate the sinusoidal functions to capture the relative position information, which is illustrated as:
@@
The output of an attention head has two terms: a term related to value (i.e., $YW_{V}$), and a term related to the relative positions. The $i$-th row of the output is calculated by
@@ 3.3.1 Image Self-Attention (ISA) Module
Figure 4: The overview of the image self-attention (ISA) module.
@@ 3.3.3 Shape Self-Attention (SSA) Module
The architecture is depicted in Fig. 5. The SSA module aims to learn contextual dependencies across the points of a shape, where individual points are associated with embedding vectors that collectively form the shape feature matrix ${X_{S}}\in{\mathbb{R}^{N_{p}\times E}}$. $N_{p}$ is the number of points/tokens of the shape and $E$ is the embedding/feature dimension. A row in the feature matrix ${X_{S}}$ is a token in the attention mechanism described in Section A. For self-attention calculation, the position of a token is defined on the undeformed template with normalized coordinates between -1 and 1. The SSA module has multiple self-attention layers to generate a sequence of feature matrices.
@@ 3.5 Error Estimation Using a Modified TransDeformer
Figure 7: The overview of the shape error estimation model (unit: mm).
@@ 4.1 Dataset
Our dataset [11] consists of de-identified lumbar spine MR images of 100 patients. Three experts annotated and reviewed each mid-sagittal MR image to ensure accuracy and consistency. Each mid-sagittal MR image was manually marked with boundaries and landmarks of the lumbar vertebrae and discs, which is guided by the established protocol [7]. The images are resized to 512 $\times$ 512. Each image is accompanied with 12 segmentation masks, corresponding to the background, 6 vertebrae (named L1, L2, L3, L3, L4, L5, and S1), and 5 discs (named D1, D2, D3, D4, and D5). We split the entire dataset into 70 training samples, 10 validation samples, and 20 test samples. A special data augmentation method was applied to the original dataset, resulting in 7000 training samples, 250 validation samples, and 2500 test samples. The data augmentation method is described in detail in our technical report [11]. Our experiments were conducted on the augmented datasets that are publicly available [11]. Data diversity is illustrated in Fig. 8 by comparing the template with the ground-truth meshes, for which the centroids are aligned.
Figure 8: Distribution of Average Point-to-Point Distance (APPD) between the template mesh and the ground truth meshes of the patients in the test set. APPD is defined in Section 4.4.
@@
4.2 Comparison with Related Approaches
4.2.1 UNet-GCN Model
This model was originally proposed by Kong et al. [22] for reconstructing human heart geometries from 3D volumetric images. The model has a UNet backbone pretrained for binary segmentation, and the encoder of the UNet is used as the image feature extractor at different scales. Three GCN modules are used to deform a template sequentially by using image features sampled at the points of the template. For our application, we followed the original design as close as possible, and made necessary adjustment by changing the 3D UNet to a 2D UNet to handle 2D sagittal images of lumbar spine. The UNet-GCN model uses GCN layers to fuse information from neighbour points for displacement prediction. Our UNet-DeformSA model has shape self-attention modules to integrate information from different points that could be far away from each other (i.e., utilizing long-range dependency), which is the key difference from and advantage over the UNet-GCN model.
4.2.2 UNet-Disp Model
4.2.3 UNet-MLP Model
4.3 Implementation
We used Pytorch ver. 2.0.0 to implement the overall frameworks, and used Pytorch Geometric ver. 2.3.0 to implement GCNs. We used the Adam optimizer with a fixed learning rate of 1e-4 and a batch size of 20. The training of each model completed within 24 hours on a single NVIDIA RTX A6000 with 48GB VRAM.
4.4 Metrics and Results of Geometry Reconstruction
To measure the accuracy of a reconstructed lumbar object (a vertebra or a disc), we assessed two key metrics: (1) Average Point-to-Point Distance (APPD) between a predicted shape and the corresponding ground-truth shape and (2) Dice Similarity Coefficient (Dice), measuring the overlap between areas enclosed by predicted and ground-truth shapes. For each metric, mean, standard deviation (std), the worst case, the $95$th or $5$th percentile among the test samples are calculated.
@@ 4.8 Hyperparameter Tuning on Validation Set
For the UNet-DeformSA model, we investigated the effect of the number of attention layers in the SSA module, and the results on the validation set are reported in Table 8. Based on this results, 2 layers of attention are used in the main experiments. For the UNet-GCN model, we evaluated different types of GCN layers (Table 9) and chose ResGatedGraphConv [28] based on the results on the validation set. For the UNet-Disp model, we explored the hyperparameter $\sigma$ (Table 10) and set it to 1 based on the results on the validation set. The backbone UNet structure is the same for the three models.
@@
Note: TransDeformer results (lower is better)
@@
Note: UNet-DeformSA results (lower is better)
Table 9: Hyperparameter Tuning on the type of GCN layers
GCN layer type APPD (mm)
mean std max Q95
ResGatedGraphConv 0.779 0.953 12.276 1.201
ChebConv 0.936 1.003 11.913 2.104
GATConv 2.191 2.993 15.890 9.587
TransformerConv 1.595 2.395 14.250 7.154
Note: UNet-GCN results (lower is better)
Table 10: Hyperparameter Tuning on the scaling factor $\sigma$
$\sigma$ APPD (mm)
mean std max Q95
1 0.817 0.188 1.363 1.149
2 0.842 0.243 2.289 1.195
3 0.899 0.366 2.660 1.531
Note: UNet-Disp results (lower is better)
@@ 5 Discussion
All of the current ML models for geometry reconstruction, including our models, are data-driven and therefore have no guarantee in output accuracy for an input image. Thus, the geometry output from a ML model has to be checked (and modified if necessary) by a human operator to ensure high accuracy for clinical use. The shape error estimation model, which is a modified version of TransDeformer, facilitates this quality control process by ranking the model-reconstructed geometries by the estimated errors. To reduce human negligence, those with larger (estimated) errors will be checked first when the operator is vigilant; and those with smaller (estimated) errors will be checked later when the operator possibly becomes fatigue.